OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (0, ) and (0, ).

OpenStudy (anonymous):

(0, 1/2) and (0, 3/4)

OpenStudy (anonymous):

midpoint = (x1 + x2)/2, (y1 + y2)/2 midpoint = (0+0)/2 , (1/2 + 3/4)/2